# A halo model approach for mock catalogs of time-variable strong gravitational lenses

### v4

@@ 3.4 Stellar mass fundamental plane
with $\sigma_{e}$ in units of $\mathrm{km/s}$ and $r_{e}$ in units of $\mathrm{kpc}$. Zahid et al. (2016, Z16 hereafter) provide the best-fit values for $(a,b,c,d)$ in Eq. (27) as $(-0.84,1.63,-1.00,4.42)$. It is known that this plane is almost invariant as a function of redshift, although the zero point may slightly evolve with redshift.
Figure 5 shows our mock galaxy sample with masses larger than $3\times 10^{10}M_{\odot}$ projected onto the MFP defined by Eq. (27) with the parameters of Z16. We find that our galaxy population is well distributed on this MFP. The root-mean-square orthogonal scatter in the MFP for our sample is $0.12~{}\mathrm{dex}$, almost independently of redshift. Fig. 7 of Z16 shows the SDSS galaxy samples at $z<0.1$ and the hCOSMOS samples at $z<0.6$ on the MFP, which are in good agreement with our result in Fig. 5. The root-mean-square orthogonal scatters in the MFP for the SDSS and hCOSMOS sample are $0.17~{}\mathrm{dex}$ and $0.24~{}\mathrm{dex}$, including not only the intrinsic scatter but also measurement errors, respectively. It should be noted that our mock galaxy population in Fig. 5 is not filtered by any observational selection function, such as the one from which the MFP relation is taken. Also, the sharp galaxy population with small $r_{e}$ on the MFP in Fig. 5 comes from the cutoff of the stellar mass to create this galaxy population.
Figure 5: Our mock galaxy catalog is plotted on the MFP. The MFP is described by Eq. (27) with parameters of $(a,b,c,d)=(-0.84,1.63,-1.00,4.42)$, which is obtained for quiescent galaxies in Hyde and Bernardi (2009). The blue dashed line displays the best-fit line for the SDSS sample of quiescent galaxies obtained in Z16.
@@ 4 Create a mock catalog of time-variable strong lenses
[…]
@@
Since the brightness of SNe changes drastically with time, the definition of the magnitude limit to detect lensed SNe is non-trivial. We follow OM10 and set the condition in which lensed SNe will be detected if their peak magnitude is more than $0.7$ mag brighter than the magnitude limit for each visit. Additionally, we consider the effective survey duration $t_{\mathrm{eff}}$ to predict the total number of lensed SNe. In the case of LSST, we assume that the effective survey period is about $3$ months each year out of 10 years of the observation, i.e., $t_{\mathrm{eff}}=2.5~{}\mathrm{yr}$, following OM10. We note that this effective survey period is a conservative estimate. In reality, a longer observation period is planned, which could result in the discovery of more lensed SNe. However, we set this effective survey period for the comparison with the OM10+ results.
@@
We largely follow OM10 for the specific procedure to generate mock catalogs, which we described below. For a given survey area and depth, we first generate a list of sources according to the adopted source luminosity function. To take account of the magnification bias, we actually adopt a 3 mag deeper limit than the magnitude limit of the survey for generating the source catalog. Each source in the sample contains information on only the redshift and apparent magnitude without any lensing magnification, and hence we ignore spatial correlations between sources. As a function of the halo mass and redshift, we also pre-compute the size of a square boundary in the source plane such that the magnification at the boundary is $\mu=1.5$ along the major axis direction, assuming an elliptical halo and galaxy with an extreme ellipticity of $0.8$, in order to be conservative. We then generate a list of halos and subhalos according the adopted halo mass function and subhalo mass function (see Sec. 2), again ignoring any spatial correlations between halos as well as between subhalos in a halo. For each halo and subhalo, we assign an ellipticity, position angle, and stellar component, adopting the model detailed in Sec. 2. For each halo and subhalo, we randomly distribute source objects from the pre-computed source catalog within a square region in the source plane whose size is pre-computed as mentioned above, assuming the Poisson distribution. For each realization of sources, we solve the lens equation using glafic, and record only events with multiple images, or with magnifications larger than some threshold even for single-image systems, which are used for the analysis in Sec. 4.4. We also check the detection criteria for the magnitude limit and the flux ratio mentioned above to construct the final sample of the mock catalog.
@@
The catalog contains all the necessary information for each mock lens, including the properties of lenses and sources, image positions, magnifications, time delays between image pairs. In practice, the catalog is five times (QSOs) or ten times (SNe) oversampled in order to reduce the shot noise. We caution that our calculations use the Salpeter IMF ($f_{\mathrm{IMF}}=1.715$) for the fiducial calculation. For comparison, mock catalogs are also produced with the Chabrier IMF ($f_{\mathrm{IMF}}=1.0$). However, the results in the Chabrier IMF case are only mentioned if the behavior differs significantly from that of the Salpeter IMF case. Additionally, we note that our mock catalog does not consider lens objects with halo masses below $10^{11}~{}M_{\odot}$ (or $10^{10}~{}M_{\odot}$ for subhalos) for a numerical reason. Therefore, it may underestimate the number of lenses with small image separation or short time delay, which will be discussed again later.
@@
where $\kappa_{\mathrm{env,hh}}$ and $\kappa_{\mathrm{env,cen}}$ are the environmental convergence due to the host halo and the central galaxy, respectively. Since the host halo follows the elliptical NFW profile, $\kappa_{\mathrm{env,hh}}$ is given by (Bartelmann, 1996)
@@
On the other hand, since the central galaxy follows the elliptical Hernquist profile, $\kappa_{\mathrm{env,cen}}$ can be written by (Keeton, 2001)
@@ 4.1 Mock catalog of lensed QSOs
First, we examine properties of mock catalogs for lensed QSOs with multiple images. The total number in the fiducial mock catalog is $4004$ for lensed QSOs, which is similar to but is somewhat smaller than the result of $4286$ from OM10+. Of these, $389$ lenses, or approximately $10\%$ of all lenses, are primarily produced by subhalos. The fraction of lensing by subhalos in our mock catalog is broadly consistent with a previous analytic estimate (Oguri, 2006). We find that the mock catalog with the Chabrier IMF instead of the Salpeter IMF includes $2169$ lensed QSOs, which is significantly smaller compared with the case assuming the fiducial Salpeter IMF. Of these, $244$ lenses are primarily produced by subhalos, indicating that the ratio is comparable to that in the fiducial mock catalog. When we add a criterion that the maximum image separation, which is defined by the maximum angular separation between any pair of multiple images, should be larger than $\Delta\theta_{\mathrm{min}}=(2/3)\theta_{\mathrm{PSF}}$, the total numbers of lensed QSOs reduces to $3471$ in the fiducial case, $1731$ in the Chabrier IMF case, and $3013$ in the case of OM10+. Since the number of lenses predicted by OM10+ has been shown to be approximately consistent with observations (see e.g., Lemon et al., 2023), our result with the fiducial Salpeter IMF should also be consistent with observations. In contrast, since the number of lenses predicted in the case of the Chabrier IMF is about half of the expectation in OM10+, our model with the Chabrier IMF apparently has difficulty in explaining the observed abundance of strongly lensed QSOs. We explicitly check the consistency with observations in Sec. 5. The above results regarding the total number of lenses are summarized in Table 3.
@@
Basically, our catalogs consist of three- or five-image lenses in accordance with the odd number theorem (see, e.g., Burke, 1981). However, in most cases, the images near the centers of lensing galaxies are strongly demagnified and unobservable. We refer to three- and five-image lenses as double (two-image) and quadruple (quad; four-image) lenses to keep the notation consistent with the one in OM10+, which basically predicts double or quad lenses from the SIE plus external shear model. We find that the fraction of quad lenses in the fiducial (Chabrier) model is approximately $10\%~{}(14\%)$, while the one in OM10+ is about $15\%$, which is slightly higher than our results. Note that these quad fractions are lower than the current observed values, about $14\%$ (see e.g., Lemon et al., 2023), which is understood from the viewpoint that the current sample of known lenses is more strongly affected by magnification bias, and hence has higher quad fractions, compared to the lenses expected from LSST. We will discuss the detail in Sec. 5.
@@
QSOs Salpeter $4004\pm 30$ ($3471\pm 26$) $403\pm 7$ $389\pm 8$ $5692\pm 94$ $570\pm 18$
QSOs Chabrier $2169\pm 22$ ($1731\pm 12$) $295\pm 18$ $244\pm 16$ $4905\pm 65$ $469\pm 25$
@@
SNe Salpeter $236\pm 15$ ($197\pm 13$) $55\pm 7$ $32\pm 7$ $861\pm 19$ $180\pm 14$
SNe Chabrier $133\pm 10$ ($100\pm 9$) $31\pm 4$ $20\pm 5$ $643\pm 23$ $125\pm 6$
Here, although the discussion in this paper is primarily tailored for future LSST observations, we explore how the number of observable strong lenses increases with the survey depth. In the left panel of Fig. 6, we plot the number of lensed QSOs over an area of $20,000~{}\mathrm{deg}^{2}$ for both our fiducial Salpeter IMF and Chabrier IMF mock catalogs as a function of the $i$-band limiting magnitude, $i_{\mathrm{lim}}$. The result for the Salpeter IMF is consistent with that of OM10 (Fig. 3), showing a shallow slope and indicating that survey area is more crucial than survey depth for discovering many lensed QSOs. The result for the Chabrier IMF is similar to that of the Salpeter IMF, but showing the smaller number of lenses. The double- and quad-fractions are also in agreement with the results of OM10, showing a gradual decrease with increasing $i_{\mathrm{lim}}$ The right panel depicts the results of lensed SNe for both our fiducial Salpeter IMF and Chabrier IMF mock catalogs over an area of $20,000~{}\mathrm{deg}^{2}$ with the effective survey period of $2.5$-yr. The slope of the number counts is steeper than the one for the lensed QSOs, suggesting that the survey depth is relatively more important for lensed SNe. The steeper dependence of the number of strong lenses on the limiting magnitudes for lensed SNe originates from lower source redshifts ($z\lesssim 1$) of lensed SNe where the gravitational lensing probability is more sensitive to the source redshift. As for OM10, we find that the fraction of quad lens systems tends to be higher than the one for the lensed QSOs, due to the very large magnification bias associated with the bright lenses, and it changes rapidly as a function of the limiting magnitude.
@@
Figure 6: The expected number of lensed QSOs (left panel) and lensed SNe (right panel) as a function of the $i$-band limiting magnitude, $i_{\mathrm{lim}}$. Here, we assume the survey area of $20,000~{}\mathrm{deg^{2}}$, and the effective survey period of $2.5$-yr for lensed SNe. The lower panels show the ratio of quad (four-image) lenses to the total number of lenses as a function of $i_{\mathrm{lim}}$. The red and green lines show the results from our fiducial Salpeter IMF mock catalogs for lensed QSOs and SNe, respectively, while the blue and purple lines represent those from the Chabrier IMF mock catalogs. Quad-lensed SNe in the Chabrier IMF mock catalogs is very rare at $i_{\mathrm{lim}}<20$, and as a result the fraction becomes 00 or $1$ in our calculation.
To check the observability of multiple images, it is essential to investigate the distribution of image separations between multiple images. Fig. 7 shows the distribution functions of maximum image separation between multiple images in our mock lenses. We find that the distribution function from our mock catalogs has a different peak from OM10+, and several mock lenses display a large image separation, $\Delta\theta>10~{}\mathrm{arcsec}$, mainly produced by galaxy groups or clusters, unlike the result of OM10+. This is one of the major results obtained by adopting the compound lens model described in Sec. 2. Note that OM10+ adopts the SIE lens model, focuses on massive galaxies as lens objects, and does not predict the lenses produced by more massive objects such as galaxy groups and clusters. For example, our fiducial model predicts that LSST 10-year observations will find about 80 lensed QSOs with the maximum image separations larger than $10~{}\mathrm{arcsec}$. We also find that the overall number of lenses in the Chabrier IMF case is lower than in the Salpeter IMF case, except for cluster-scale lenses with maximum image separations larger than 10 arcsec for which dark matter rather than the stellar mass dominates the lens potential. Furthermore, there are notable differences compared with OM10+, even on the small image separation side. A caveat is that since we set a lower limit on the mass of halos to be calculated as $M_{\mathrm{hh}}>10^{11}~{}M_{\odot}$ for a numerical reason, we may underestimate lenses with small image separations. We confirm that the number of lenses with image separations of $\Delta\theta\geq(2/3)\theta_{\mathrm{PSF}}$ is not affected by this lower limit on the halo mass.
@@
An advantage of mock catalogs is that we can easily check how distributions of lens and source parameters differ from those of unlensed populations. As specific examples, Fig. 9 shows probability distributions of the lens ellipticity and the alignment between the lens galaxy and lens halo separately for double and quad lenses. We find that the lensing halos and galaxies of quad lens systems typically have higher ellipticity than those of double lenses, although not as pronounced as in quad lens systems in OM10+. Our result is broadly consistent with previous studies (see, e.g., Rozo et al., 2007).
@@
The probability distributions of the position angle of lens halos and galaxies against the position angle of external shear are shown in the left and middle panels in Fig. 10. We also plot the distribution of external shear itself in the right panel of Fig. 10. It appears that the probability distribution in quad lenses shows slight tails toward the higher values compared to the one in double lenses. We conduct the Kolmogorov-Smirnov test at $\gamma_{\mathrm{ext}}$ to confirm that the difference of the tails is significant with the p-value of $5\times 10^{-4}$. We also find that the relative orientation of lens halos and galaxies against external shear is biased for both cases of double and quad lenses, albeit weakly, compared to OM10+. That is, for double lenses, the direction of external shear is more likely to be aligned with the minor axis of the lens halo and lens galaxy, whereas, for quad lenses, the direction of external shear is more likely to be aligned with the major axis. This bias can be understood through the degeneracy between the ellipticity of the lens object and the external shear (see, e.g., Keeton et al., 1997). However, our mock catalogs exhibit a more complex bias than OM10+ due to the compound lens model as described in Sec. 2. It may be related with their differences in the distributions of lens ellipticities discussed in Fig. 9.
@@
One of the quantities that are related to the maximum image separation is the velocity dispersion of galaxies, which is displayed in Fig. 11 in the same manner as in Fig. 8. We remind the reader that OM10+ adopts the velocity dispersion function for $\sigma_{e/8}$ based on Oguri (2018), which is shown in Eq. (23) with the redshift evolution in Eq. (24). We convert $\sigma_{e/8}$ to $\sigma_{e}$ using Eq. (25) to fairly compare both results. The distributions of maximum image separation and velocity dispersion show similar shapes because of their tight relation, e.g., $\Delta\theta\propto\sigma_{e}^{2}$ in the SIS case. As noted above, the reason why lenses with larger $\sigma_{e}$ are expected compared to OM10+ is that our lens model is able to represent galaxy groups and galaxy clusters. In addition, we may underestimate the number of lenses at small velocity dispersions due to the cutoff of small-mass halos.
@@
$\begin{split}M_{\text{lh}}=\left\{\begin{array}[]{lr}M_{\mathrm{f}}&\text{ (lenses by subhalos) }\\ M_{\mathrm{hh}}&\text{ (otherwise) },\end{array}\right.\end{split}$ (32)
@@
$\begin{split}M_{\text{l*}}=\left\{\begin{array}[]{lr}M_{\mathrm{s*}}&\text{ (lenses by subhalos) }\\ M_{\mathrm{c*}}&\text{ (otherwise) },\end{array}\right.\end{split}$ (33)

### v5

@@ 3.4 Stellar mass fundamental plane
with $\sigma_{e}$ in units of $\mathrm{km/s}$ and $r_{e}$ in units of $\mathrm{kpc}$. 2016ApJ...821..101Z provide the best-fit values for $(a,b,c,d)$ in Eq. (27) as (0.84,1.63,1.00,4.42)(-0.84,1.63,-1.00,4.42). It is known that this plane is almost invariant as a function of redshift, although the zero point may slightly evolve with redshift.
Figure 5 shows our mock galaxy sample with masses larger than $3\times 10^{10}M_{\odot}$ projected onto the MFP defined by Eq. (27) with the parameters of 2016ApJ...821..101Z. We find that our galaxy population is well distributed on this MFP. The root-mean-square orthogonal scatter in the MFP for our sample is $0.12~\mathrm{dex}$, almost independently of redshift. Fig. 7 of 2016ApJ...821..101Z shows the SDSS galaxy samples at $z<0.1$ and the hCOSMOS samples at $z<0.6$ on the MFP, which are in good agreement with our result in Fig. 5. The root-mean-square orthogonal scatters in the MFP for the SDSS and hCOSMOS sample are $0.17~\mathrm{dex}$ and $0.24~\mathrm{dex}$, including not only the intrinsic scatter but also measurement errors, respectively. It should be noted that our mock galaxy population in Fig. 5 is not filtered by any observational selection function, such as the one from which the MFP relation is taken. Also, the sharp galaxy population with small $r_{e}$ on the MFP in Fig. 5 comes from the cutoff of the stellar mass to create this galaxy population.
Figure 5: Our mock galaxy catalog is plotted on the MFP. The MFP is described by Eq. (27) with parameters of (a,b,c,d)=(0.84,1.63,1.00,4.42)(a,b,c,d)=(-0.84,1.63,-1.00,4.42), which is obtained for quiescent galaxies in 2009MNRAS.396.1171H. The blue dashed line displays the best-fit line for the SDSS sample of quiescent galaxies obtained in 2016ApJ...821..101Z.
@@ 4 Create a mock catalog of time-variable strong lenses
As a specific example, this paper presents a mock catalog of lensed QSOs and SNe generated using our model, tailored for the baseline survey planned by LSST. We adopt the same source models of QSOs and SNe as OM10+, and update the deflector model that is detailed in Sec. 2. As in 2010MNRAS.405.2579O, we consider five types of SNe, Ia, Ib/c, IIP, IIL, and IIn. Therefore, any difference between results in this paper and those in OM10+ should originate from the difference of their lens models rather than the source models. We set criteria specific to LSST, where the survey limiting magnitude in the $i$-band is $m_{\mathrm{i,lim}}=23.3$, and the flux ratio between the brightest and the second brightest images for double lenses should be $f_{\mathrm{lim}}<0.1$ [^2] Changing the flux ratio limit from $0.1$ to $0.01$ increases the number of lensed events by approximately $12\%$.. This is because asymmetric double images with large flux differences can be very challenging to detect due to dynamic range problems produced by the brighter image as well as the obscuration of the fainter images by lensing galaxies. Following 2010MNRAS.405.2579O, we use the magnification of the fainter image (excluding the central image) for double lens systems and that of the third brightest image for the other multiple image systems for computing the magnification bias.
2010MNRAS.405.2579O sets a limit on image separation as $\Delta\theta_{\mathrm{min}}=(2/3)\theta_{\mathrm{PSF}}$, where $\theta_{\mathrm{PSF}}$ is the typical seeing full width at half-maximum (FWHM) for a given survey (e.g., $\theta_{\mathrm{PSF}}=0^{\prime\prime}.75$ for LSST). On the other hand, several approaches to search for small-image separation strong lenses with apparently unresolved multiple images have been proposed and demonstrated (2014Sci...344..396Q; 2017Sci...356..291G; 2017ApJ...834L...5G; 2023NatAs...7.1098G; 2021NatAs...5..569S; 2024MNRAS.531.3509A; 2024arXiv240415389B). Therefore, in this work, we do not impose such a lower limit on the image separation when creating the mock catalog, as adopted in OM10+.
@@
Since the brightness of SNe changes drastically with time, the definition of the magnitude limit to detect lensed SNe is non-trivial. We follow 2010MNRAS.405.2579O and set the condition in which lensed SNe will be detected if their peak magnitude is more than $0.7$ mag brighter than the magnitude limit for each visit. Additionally, we consider the effective survey duration $t_{\mathrm{eff}}$ to predict the total number of lensed SNe. In the case of LSST, we assume that the effective survey period is about $3$ months each year out of 10 years of the observation, i.e., $t_{\mathrm{eff}}=2.5~\mathrm{yr}$, following 2010MNRAS.405.2579O. We note that this effective survey period is a conservative estimate. In reality, a longer observation period is planned, which could result in the discovery of more lensed SNe. However, we set this effective survey period for the comparison with the OM10+ results.
@@
We largely follow 2010MNRAS.405.2579O for the specific procedure to generate mock catalogs, which we described below. For a given survey area and depth, we first generate a list of sources according to the adopted source luminosity function. To take account of the magnification bias, we actually adopt a 3 mag deeper limit than the magnitude limit of the survey for generating the source catalog. Each source in the sample contains information on only the redshift and apparent magnitude without any lensing magnification, and hence we ignore spatial correlations between sources. As a function of the halo mass and redshift, we also pre-compute the size of a square boundary in the source plane such that the magnification at the boundary is $\mu=1.5$ along the major axis direction, assuming an elliptical halo and galaxy with an extreme ellipticity of $0.8$, in order to be conservative. We then generate a list of halos and subhalos according the adopted halo mass function and subhalo mass function (see Sec. 2), again ignoring any spatial correlations between halos as well as between subhalos in a halo. For each halo and subhalo, we assign an ellipticity, position angle, and stellar component, adopting the model detailed in Sec. 2. For each halo and subhalo, we randomly distribute source objects from the pre-computed source catalog within a square region in the source plane whose size is pre-computed as mentioned above, assuming the Poisson distribution. For each realization of sources, we solve the lens equation using glafic, and record only events with multiple images, or with magnifications larger than some threshold even for single-image systems, which are used for the analysis in Sec. 4.4. We also check the detection criteria for the magnitude limit and the flux ratio mentioned above to construct the final sample of the mock catalog.
@@
The catalog contains all the necessary information for each mock lens, including the properties of lenses and sources, image positions, magnifications, time delays between image pairs. In practice, the catalog is five times (QSOs) or ten times (SNe) oversampled in order to reduce the shot noise. We caution that our calculations use the Salpeter IMF ($f_{\mathrm{IMF}}=1.715$) for the fiducial calculation. For comparison, mock catalogs are also produced with the Chabrier IMF ($f_{\mathrm{IMF}}=1.0$). However, the results in the Chabrier IMF case are only mentioned if the behavior differs significantly from that of the Salpeter IMF case. Additionally, we note that our mock catalog does not consider lens objects with halo masses below $10^{10}~M_{\odot}$ (or $10^{9}~M_{\odot}$ for subhalos) for a numerical reason. Therefore, it may underestimate the number of lenses with small image separation or short time delay, which will be discussed again later.
@@
where $\kappa_{\mathrm{env,hh}}$ and $\kappa_{\mathrm{env,cen}}$ are the environmental convergence due to the host halo and the central galaxy, respectively. Since the host halo follows the elliptical NFW profile, $\kappa_{\mathrm{env,hh}}$ is given by (1996A&A...313..697B)
@@
On the other hand, since the central galaxy follows the elliptical Hernquist profile, $\kappa_{\mathrm{env,cen}}$ can be written by (2001astro.ph..2341K)
@@ 4.1 Mock catalog of lensed QSOs
First, we examine properties of mock catalogs for lensed QSOs with multiple images. The total number in the fiducial mock catalog is $4907$ for lensed QSOs, which is similar to but is somewhat larger than the result of $4286$ from OM10+. Of these, $1297$ lenses, or approximately $25\%$ of all lenses, are primarily produced by subhalos. The fraction of lensing by subhalos in our mock catalog is broadly consistent with a previous analytic estimate (2006MNRAS.367.1241O). We find that the mock catalog with the Chabrier IMF instead of the Salpeter IMF includes $2732$ lensed QSOs, which is significantly smaller compared with the case assuming the fiducial Salpeter IMF. Of these, $800$ lenses are primarily produced by subhalos, indicating that the ratio is comparable to that in the fiducial mock catalog. When we add a criterion that the maximum image separation, which is defined by the maximum angular separation between any pair of multiple images, should be larger than $\Delta\theta_{\mathrm{min}}=(2/3)\theta_{\mathrm{PSF}}$, the total numbers of lensed QSOs reduces to $4223$ in the fiducial case, $2179$ in the Chabrier IMF case, and $3013$ in the case of OM10+. Since the number of lenses predicted by OM10+ has been shown to be approximately consistent with observations (see e.g., 2023MNRAS.520.3305L), our result should also be consistent with observations. We explicitly check the consistency with observations in Sec. 5. The above results regarding the total number of lenses are summarized in Table 3.
@@
Basically, our catalogs consist of three- or five-image lenses in accordance with the odd number theorem (see, e.g., 1981ApJ...244L...1B). However, in most cases, the images near the centers of lensing galaxies are strongly demagnified and unobservable. We refer to three- and five-image lenses as double (two-image) and quadruple (quad; four-image) lenses to keep the notation consistent with the one in OM10+, which basically predicts double or quad lenses from the SIE plus external shear model. We find that the fraction of quad lenses in the fiducial (Chabrier) model is approximately $10\%$ ($13\%$), while the one in OM10+ is about $15\%$, which is slightly higher than our results. Note that these quad fractions are lower than the current observed values, about $14\%$ (see e.g., 2023MNRAS.520.3305L), which is understood from the viewpoint that the current sample of known lenses is more strongly affected by magnification bias, and hence has higher quad fractions, compared to the lenses expected from LSST. We will discuss the detail in Sec. 5.
@@
QSOs Salpeter $4907\pm 53$ ($4223\pm 62$) $514\pm 25$ $1297\pm 29$ $7459\pm 96$ $764\pm 19$
QSOs Chabrier $2732\pm 43$ ($2179\pm 34$) $362\pm 19$ $800\pm 25$ $6422\pm 67$ $629\pm 19$
@@
SNe Salpeter $288\pm 17$ ($234\pm 12$) $69\pm 7$ $85\pm 10$ $1098\pm 19$ $230\pm 8$
SNe Chabrier $169\pm 18$ ($127\pm 15$) $39\pm 6$ $54\pm 9$ $822\pm 30$ $170\pm 11$
Here, although the discussion in this paper is primarily tailored for future LSST observations, we explore how the number of observable strong lenses increases with the survey depth. In the left panel of Fig. 6, we plot the number of lensed QSOs over an area of $20,000~\mathrm{deg}^{2}$ for both our fiducial Salpeter IMF and Chabrier IMF mock catalogs as a function of the $i$-band limiting magnitude, $i_{\mathrm{lim}}$. The result for the Salpeter IMF is consistent with that of 2010MNRAS.405.2579O (Fig. 3), showing a shallow slope and indicating that survey area is more crucial than survey depth for discovering many lensed QSOs. The result for the Chabrier IMF is similar to that of the Salpeter IMF, but showing the smaller number of lenses. The double- and quad-fractions are also in agreement with the results of 2010MNRAS.405.2579O, showing a gradual decrease with increasing $i_{\mathrm{lim}}$ The right panel depicts the results of lensed SNe for both our fiducial Salpeter IMF and Chabrier IMF mock catalogs over an area of $20,000~\mathrm{deg}^{2}$ with the effective survey period of $2.5$-yr. The slope of the number counts is steeper than the one for the lensed QSOs, suggesting that the survey depth is relatively more important for lensed SNe. The steeper dependence of the number of strong lenses on the limiting magnitudes for lensed SNe originates from lower source redshifts ($z\lesssim 1$) of lensed SNe where the gravitational lensing probability is more sensitive to the source redshift. As for 2010MNRAS.405.2579O, we find that the fraction of quad lens systems tends to be higher than the one for the lensed QSOs, due to the very large magnification bias associated with the bright lenses, and it changes rapidly as a function of the limiting magnitude.
@@
Figure 6: The expected number of lensed QSOs (left panel) and lensed SNe (right panel) as a function of the $i$-band limiting magnitude, $i_{\mathrm{lim}}$. Here, we assume the survey area of $20,000~\mathrm{deg^{2}}$, and the effective survey period of $2.5$-yr for lensed SNe. The lower panels show the ratio of quad (four-image) lenses to the total number of lenses as a function of $i_{\mathrm{lim}}$. The red and green lines show the results from our fiducial Salpeter IMF mock catalogs for lensed QSOs and SNe, respectively, while the blue and purple lines represent those from the Chabrier IMF mock catalogs. Quad-lensed SNe in the Chabrier IMF mock catalogs is very rare at $i_{\mathrm{lim}}<20$, and as a result the fraction becomes 0 or 11 in our calculation.
To check the observability of multiple images, it is essential to investigate the distribution of image separations between multiple images. Fig. 7 shows the distribution functions of maximum image separation between multiple images in our mock lenses. We find that the distribution function from our mock catalogs has a different peak from OM10+, and several mock lenses display a large image separation, $\Delta\theta>10~\mathrm{arcsec}$, mainly produced by galaxy groups or clusters, unlike the result of OM10+. This is one of the major results obtained by adopting the compound lens model described in Sec. 2. Note that OM10+ adopts the SIE lens model, focuses on massive galaxies as lens objects, and does not predict the lenses produced by more massive objects such as galaxy groups and clusters. For example, our fiducial model predicts that LSST 10-year observations will find about 80 lensed QSOs with the maximum image separations larger than $10~\mathrm{arcsec}$. We also find that the overall number of lenses in the Chabrier IMF case is lower than in the Salpeter IMF case, except for cluster-scale lenses with maximum image separations larger than 10 arcsec for which dark matter rather than the stellar mass dominates the lens potential. Furthermore, there are notable differences compared with OM10+, even on the small image separation side. A caveat is that since we set a lower limit on the mass of halos to be calculated as $M_{\mathrm{hh}}>10^{10}~M_{\odot}$ for a numerical reason, we may underestimate lenses with small image separations. We confirm that the number of lenses with image separations of $\Delta\theta\geq(2/3)\theta_{\mathrm{PSF}}$ is not affected by this lower limit on the halo mass.
@@
An advantage of mock catalogs is that we can easily check how distributions of lens and source parameters differ from those of unlensed populations. As specific examples, Fig. 9 shows probability distributions of the lens ellipticity and the alignment between the lens galaxy and lens halo separately for double and quad lenses. We find that the lensing halos and galaxies of quad lens systems typically have higher ellipticity than those of double lenses, although not as pronounced as in quad lens systems in OM10+. Our result is broadly consistent with previous studies (see, e.g., Rozo:2007bj).
@@
The probability distributions of the position angle of lens halos and galaxies against the position angle of external shear are shown in the left and middle panels in Fig. 10. We also plot the distribution of external shear itself in the right panel of Fig. 10. It appears that the probability distribution in quad lenses shows slight tails toward the higher values compared to the one in double lenses. We conduct the Kolmogorov-Smirnov test at $\gamma_{\mathrm{ext}}$ to confirm that the difference of the tails is significant with the p-value of $3\times 10^{-10}$. We also find that the relative orientation of lens halos and galaxies against external shear is biased for both cases of double and quad lenses, albeit weakly, compared to OM10+. That is, for double lenses, the direction of external shear is more likely to be aligned with the minor axis of the lens halo and lens galaxy, whereas, for quad lenses, the direction of external shear is more likely to be aligned with the major axis. This bias can be understood through the degeneracy between the ellipticity of the lens object and the external shear (see, e.g., 1997ApJ...482..604K). However, our mock catalogs exhibit a more complex bias than OM10+ due to the compound lens model as described in Sec. 2. It may be related with their differences in the distributions of lens ellipticities discussed in Fig. 9.
@@
One of the quantities that are related to the maximum image separation is the velocity dispersion of galaxies, which is displayed in Fig. 11 in the same manner as in Fig. 8. We remind the reader that OM10+ adopts the velocity dispersion function for $\sigma_{e/8}$ based on 2018MNRAS.480.3842O, which is shown in Eq. (23) with the redshift evolution in Eq. (24). We convert $\sigma_{e/8}$ to $\sigma_{e}$ using Eq. (25) to fairly compare both results. The distributions of maximum image separation and velocity dispersion show similar shapes because of their tight relation, e.g., $\Delta\theta\propto\sigma_{e}^{2}$ in the SIS case. As noted above, the reason why lenses with larger $\sigma_{e}$ are expected compared to OM10+ is that our lens model is able to represent galaxy groups and galaxy clusters. In addition, we may underestimate the number of lenses at small velocity dispersions due to the cutoff of small-mass halos.
@@
[…]
@@
[…]